# Revealing the nature of ultra-long period objects with space-based gravitational-wave interferometers

### v2

@@ Abstract
A few members of the recently-discovered class of ultra-long period objects have been identified as binaries with white-dwarf primaries. In most cases however, electromagnetic data are inconclusive and isolated magnetars or compact binaries remain viable. If the pulsation period matches that of the orbit though – as is the case for ILT J1101+5521 and GLEAM-X J0704–37 – some of these elusive radio transients could be gravitational-wave bright in the mHz band. Space-based interferometers could thus be used to provide independent constraints on their nature. We quantify the signal-to-noise ratio for the known systems, under a variety of scenarios, and show that a few could be detectable for sufficiently large chirp masses. Astrophysical implications for (non-)detections are discussed.
@@ 1 Introduction
Ultra-long period objects (ULPs) are an elusive class of coherent, radio-pulsating transients with periods of $P\sim$ hours that have thus far evaded classification. Their brightness variations imply compact sources and, given spectral similarities to pulsar emissions, the natural interpretation is that they are slow neutron stars (see Coti Zelati & Borghese, 2024, for a discussion). However, this is theoretically challenging to accept as their radio luminosities ($L_{\nu}$) categorically exceed the implied spindown luminosity, necessitating an alternative power reservoir. While magnetic fields could provide such a battery for magnetar progenitors (Beniamini, Wadiasingh & Metzger, 2020; Rea et al., 2022; Cooper & Wadiasingh, 2024), optical spectroscopy has confirmed that some ULPs are binaries consisting of white dwarfs with M-dwarf companions (WDMDs; Hurley-Walker et al., 2024; de Ruiter et al., 2025). As observational efforts are revealing a sizeable population of ULPs, residing within new portions of the $P$-$\dot{P}$, $L_{\nu}$-$L_{\rm X}$, and other parameter spaces, tools to constrain their nature are of astrophysical relevance.
@@
Because pulsations from ILT J1101+5521 and GLEAM-X J0704–37 – confirmed WDMDs – are phase-aligned with their orbital motion (de Ruiter et al., 2025; Rodriguez, 2025), this may be the case for other ULPs. On the other hand, the lowest-period systems cannot be WDMDs with $P\approx P_{\text{orb}}$ because the companion would not fit inside its Roche lobe. Even so, because the maser-like radio emission mechanisms often invoked to explain pulsations in binaries do not rely on companion makeup (Melrose, 2017; Qu & Zhang, 2025) a pair of compact objects, interacting magnetically in a tight orbit, are also viable progenitors (Lyman et al., 2025).
@@
We point out that because several ULPs reside at sub-kpc distances, they may be strong sources of gravitational waves (GWs) independently of the radio mechanism if the orbital period corresponds to the pulsation period. As eccentricity should be erased before a binary compactifies to $P_{\text{orb}}\sim$ hours (Peters, 1964), emissions at frequencies of $f_{\rm GW}=2/P_{\text{orb}}\approx 0.6\,\text{mHz}\times\left({1\,\text{hr}}/{P_{\text{orb}}}\right)$ ought to dominate the spectrum in this scenario. Within the next decade, a network of space-based, GW interferometers will become operational, namely the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017) plus the two Chinese-led projects Taiji (Luo et al., 2020) and TianQin (Huang et al., 2020). These detectors will be most sensitive in the $\gtrsim$ mHz band and can search for signals from ULPs. Consequently, future GW observations can be used to effectively confirm or rule out binary scenarios with two heavy objects in several systems where electromagnetic data are inconclusive.
@@
This short paper is organised as follows. Multiwavelength observations and binary scenarios more generally are reviewed in Section 2 to pave the way for estimates of the characteristic strains from ULPs in Section 3. Signal-to-noise ratios (SNRs) are quantified in Section 3.1, where it shown that several ULPs are promising candidates for GW detection under favourable conditions. Some discussion on astrophysical implications are also provided, with conclusions given in Section 4.
@@ 2 Constraints on binary nature
Given their observed brightness temperatures, compact sources are required for ULPs to emit coherently (e.g. Qu & Zhang, 2025). With this in mind, the basic question we pose is that if ULPs are (i) composed of a binary involving at least one compact object and (ii) pulsating at the orbital period, would they be visible to space-based GW interferometers? From a detection standpoint, the main ingredients needed to assess this are the distance, $d$, and the chirp mass,
$\mathcal{M}=\frac{\left(M_{\star}M_{\rm c}\right)^{3/5}}{\left(M_{\star}+M_{\rm c}\right)^{1/5}},$ (1)
for primary and companion masses $M_{\star}$ and $M_{\rm c}$, respectively. While ULP distances are reasonably constrained thanks to their dispersion measures, system masses (or indeed binary nature) are much more uncertain in most cases. In order to estimate the viable chirp-mass range and the viability of the hypothesis generally, we begin by reviewing the main constraints in this context.
@@ 2.1 White-dwarf plus M-dwarf
Two ULPs—ILT J1101 and GX J0740—have been confirmed as WDMDs with $P=P_{\rm orb}$ from broadband photometry, optical spectroscopy, and other channels (Hurley-Walker et al., 2024; de Ruiter et al., 2025; Rodriguez, 2025). Their data respectively indicate $M_{\rm c}\approx 0.19M_{\odot}$ and $M_{\rm c}\approx 0.32M_{\odot}$ (though cf. Rodriguez, 2025, for a Monte Carlo analysis). While primary masses cannot be dynamically-inferred with certainty due to the unknown inclination, the discovery papers suggest $M_{\star}\sim 0.6M_{\odot}$ and $M_{\star}\sim 0.8M_{\odot}$, respectively.
For ILT J1101 however, de Ruiter et al. (2025) note that arguments based on Roche-lobe overflow indicate the tighter bound [^1] This restriction applies for an M-dwarf radius of $R_{\rm MD}=0.217R_{\odot}$. If the radius was lower by even a few percent ($R_{\rm WD}\lesssim 0.21R_{\odot}$), the primary mass becomes practically unconstrained (see their extended data figure 8). $M_{\star}\lesssim 0.3M_{\odot}$. In particular, observations of M dwarfs indicate a mass-radius relationship of $M_{\rm MD}/M_{\odot}\approx R_{\rm MD}/R_{\odot}$ (Parsons et al., 2018), with the lightest (confirmed) object being EBLM J0555–57AB with $M_{\rm MD}=0.084(4)M_{\odot}$ (von Boetticher et al., 2019). Using the Eggleton (1983) formula and assuming a Keplerian orbit, the radius of the Roche sphere reads
@@
$\gtrsim 0.1R_{\odot}\left(\frac{P_{\rm orb}}{60\text{ min}}\right)^{2/3},$
for orbital separation $a$ and mass-ratio $q$. Since $R_{\rm MD}\leq R_{\rm R}$ is required for stability, an effective upper limit to $M_{\star}$ via $q$ is set by expression (2) if $P=P_{\rm orb}$. Such an argument further implies that it is unlikely that ULPs with $P\lesssim 30$ min are WDMDs pulsing at the orbital period (see Table 1 for our considered sample). Alternative scenarios with either more compact progenitors, $P\neq P_{\rm orb}$, or isolated stars are required in these cases.
Table 1: Pulsation periods (in descending order) and dispersion-measure distances for the ULPs considered in this work. Some source names have been shortened.
@@
Notes. ${}^{\text{a}}$ Dong et al. (2024) ${}^{\text{b}}$ Hurley-Walker et al. (2022) ${}^{\text{c}}$ Hurley-Walker et al. (2023) ${}^{\text{d}}$ Wang et al. (2025) ${}^{\text{e}}$ Caleb et al. (2024) ${}^{\text{f}}$ Hyman et al. (2009) ${}^{\text{g}}$ de Ruiter et al. (2025) ${}^{\text{h}}$ Hurley-Walker et al. (2024) ${}^{\text{i}}$ Lee et al. (2025).
With respect to the possibility that $P\neq P_{\rm orb}$, there are two known white-dwarf pulsars (J1912 and Ar Sco) with periods that do not correspond to the orbit but rather the spin of the primary (Marsh et al., 2016; Pelisoli et al., 2023, 2024). However, the radio profiles of these objects are notably different from the two confirmed WDMD ULPs. Their observed (highest peak) radio luminosities, $\sim 2\times 10^{26}$ erg/s at $\approx 1$ GHz, are much lower than the values of ILT J1101 ($\sim 1.5\times 10^{28}$ erg/s at $\approx$1 GHz) and GX J0740 ($\sim 10^{28}$ erg/s at $\approx$100 MHz); see appendix B.3 of Rodriguez (2025). Additionally, the degree of linear polarization in J1912 and Ar Sco are at the level of a few percent while the ULPs ILT J1101 and GX J0740 display values of at least $\approx 20\%$ (see table 1 in Qu & Zhang, 2025). Ar Sco and J1912 also shine persistently in X-rays, while ULPs do not. The operative pulsation mechanisms and binary characteristics are likely different therefore between these astronomical classes; we exclude J1912 and Ar Sco from our sample for these reasons.
@@ 2.2 White dwarfs with compact companions
If the shortest period ULPs are not WDMDs, what type of systems could be they be? While isolated neutron stars are viable – and perhaps even favoured in most cases (Beniamini, Wadiasingh & Metzger, 2020; Cooper & Wadiasingh, 2024) – double degenerates are harder to rule out owing to their intrinsic dimness (see Lyman et al., 2025, for a discussion).
@@ 2.2.1 CHIME J0630+25
$\dot{P}_{\rm GW}\approx-5\times 10^{-11}\left(\frac{P_{\rm orb}}{421\text{ s}}\right)^{-\tfrac{5}{3}}\left(\frac{\mathcal{M}}{0.5M_{\odot}}\right)^{\tfrac{5}{3}}\text{ss}^{-1}.$ (5)
Secondly, the lack of an optical counterpart is concerning owing to the closeness of the source ($d\sim 100$ pc), as many WDWDs are brighter (e.g. ZTF J153932.16+502738.8 with an orbital period of $\sim 7$ min is optically visible despite its $\sim 2.3$ kpc distance; Burdge et al., 2019). Dust obscuration or binary inclination could alleviate this tension to some degree, though the above pose a challenge for a WDWD interpretation. On the other hand, the fact that the best-fit $\dot{P}$ is negative hints at a binary, as emphasised by Dong et al. (2024).
All considered, we fix $\mathcal{M}_{\textrm{J}0630}=0.1M_{\odot}$ to avoid tension with expression (5), remarking that orbital decay could be accelerated by tides or electromagnetic interactions (decreasing the permitted $\mathcal{M}$) or decelerated if dissipative torques transfer angular momentum from the primary to the orbit (increasing it; see Marsh, Nelemans & Steeghs, 2004). We emphasise however that this source, together with DA J1832 (see Wang et al., 2025; Suvorov, Dehman & Pons, 2025), is probably the least likely candidate for a binary ULP. Nevertheless, the point we make in this work is that LISA can constrain binary scenarios in a way that electromagnetic data cannot: even if $\mathcal{M}$ was lower by an order of magnitude, the closeness of this source together with its short period would permit detection if $P=P_{\rm orb}$ (Sec. 3.1).
@@ 2.2.2 GX J1627 and GPM J1839
GPM J1839 stands out amongst ULPs as this source has displayed a high duty cycle ($\sim 25\%$) going back to 1988 (Hurley-Walker et al., 2023). In a neutron-star scenario, this points towards aggressive crustal activity despite being old and ‘cold’ (Cooper & Wadiasingh, 2024), which could be used to implicitly constrain magnetar evolution (e.g. Dehman et al., 2020). This is supported by observations of linear-to-circular polarisation conversion taking place in a manner remarkably reminiscent of the transient magnetar XTE J1810–197 (Men et al., 2025). If instead confirmed as a binary, independently of whether $P=P_{\rm orb}$, the fact that the radio luminosity exceeds that of AR Sco by several orders of magnitude indicates that pulsations from binaries involving dwarfs could be highly variable, important to account for in future radio surveys. Dedicated X-ray monitoring would also be useful to constrain source nature as the current upper limit, $L^{\rm max}_{\rm X}\sim 10^{33}$ erg/s, is relatively weak.
@@ 2.3 Neutron stars plus one
Figure 1: Characteristic strains, equation (8), for ULPs listed in Tab. 1 (see legends) assuming $P=P_{\text{orb}}$ and $T=2$ yr. Objects are colour-coded according to different scenarios corresponding to chirp masses $\mathcal{M}$ set by expressions (3) (green), (4) (red), or (6) (blue). Two exceptions are CHIME J0630 (Sec. 2.2.1) and GPM J1839 (Sec. 2.2.2), which could be double-compact systems but where orbital braking maxima restrict the component masses if there is no active mass transfer. Filled symbols correspond to cases which exceed or overlap with the LISA strain curve, taken from data provided by Robson, Cornish & Liu (2019).
Except where the above exclusions apply, the chirp mass could be considerably higher if invoking a compact neutron-star/white-dwarf binary like PSR J1141–6545 ($P_{\rm orb}\approx 4.7$ hr and $\mathcal{M}\sim 1\,M_{\odot}$; Bailes et al., 2003) or a neutron/helium star binary like PSR J1928+1815 ($P_{\rm orb}\approx 3.6$ hr and $\mathcal{M}\sim 1.2M_{\odot}$; Yang et al., 2025). Binaries involving mature neutron stars are particularly difficult to constrain with multiwavelength data, much like isolated objects. While it is more plausible that the pulsation periods correspond to the rotation rate of the primary in such cases – and indeed it has been argued that mass-loaded winds, magnetospheric twists, or interactions with a fallback disk could allow magnetars to reach ultra-long periods (Beniamini, Wadiasingh & Metzger, 2020; Beniamini et al., 2023; Suvorov, Dehman & Pons, 2025) – the assumption $P=P_{\rm orb}$ is difficult to exclude a priori. The final (most optimistic) scenario we consider corresponds to a J1141-like chirp mass of
@@
where the subscript indicates a degree of agnosticism with respect to a compact companion. For example, Turolla, Possenti & Treves (2005) suggested that GCRT J1745 could contain a double neutron-star system with $\mathcal{M}\approx 1.13M_{\odot}$ and $P=P_{\rm orb}$ where radio emission originates from shocks within the interaction zone of the primary’s wind and the companion’s magnetosphere.
We close this section by stressing that we have not attempted to present a complete survey of ULP scenarios, as these can be found in the respective discovery papers and literature (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). Indeed, our goal here concerns GW visibility, for which the exact binary makeup is unimportant (though is, of course, important with respect to implications of detections). Even if certain types of binarity may be disfavoured in individual cases, we consider a range of chirp masses to estimate SNRs. Larger or smaller $\mathcal{M}$ values can be easily substituted using the estimates provided in the following sections.

### v4

# Revealing the nature of long-period transients with space-based gravitational-wave interferometers

@@ Abstract
A few members of the recently discovered class of long-period transients have been identified as binaries with white-dwarf primaries. In most cases, however, electromagnetic data are inconclusive, and isolated magnetars or compact binaries remain viable. If the pulsation period matches that of the orbit—as is the case for ILT J1101+5521 and GLEAM-X J0704–37—some of these elusive radio transients could be gravitational-wave bright in the mHz band. Space-based interferometers could thus be used to provide independent constraints on their nature. We quantify the signal-to-noise ratio for the known systems under various scenarios and show that a few could be detectable for sufficiently large chirp masses. Astrophysical implications for (non)detections are discussed.
@@ 1 Introduction
Radio surveys are revealing a sizeable population of what are called ‘long period transients’ (LPTs). These systems pulse coherently in the radio band with periods that range from several minutes to many hours and exhibit diverse observational properties. LPT brightness variations imply compact sources and, given spectral similarities to pulsar emissions, the natural interpretation is that they are slowly rotating neutron stars (see Coti Zelati & Borghese, 2024, for a discussion). However, this is theoretically challenging to accept as their reported radio luminosities ($L_{\nu}$) typically exceed the implied spindown luminosity (e.g. Hurley-Walker et al., 2022; Men et al., 2025), necessitating an alternative power reservoir. While magnetic fields could provide such a battery for magnetar progenitors (Beniamini, Wadiasingh & Metzger, 2020; Cooper & Wadiasingh, 2024), optical spectroscopy has confirmed that some LPTs are binaries consisting of white dwarfs with M-dwarf companions (WDMDs; Hurley-Walker et al., 2024; de Ruiter et al., 2025; Rodriguez, 2025). As LPTs reside within largely unexplored parameter spaces, tools to constrain their nature are of astrophysical relevance.
@@
Because pulsations from ILT J1101+5521 and GLEAM-X J0704–37 – confirmed WDMDs – are phase-aligned with their orbital motion (de Ruiter et al., 2025; Rodriguez, 2025), this may also apply to other LPTs. However, the lowest-period systems cannot be WDMDs with pulsation periods locked to the orbit ($P\approx P_{\text{orb}}$), as the companion would not fit within its Roche lobe. Nevertheless, since the maser-like radio emission mechanisms often invoked to explain pulsations in binaries do not rely on companion makeup (Melrose, 2017; Qu & Zhang, 2025), LPTs could also involve a pair of compact objects (Dong et al., 2025b; Bloot et al., 2025; Lyman et al., 2025).
@@
We point out that because several LPTs reside at sub-kpc distances, they may be strong sources of gravitational waves (GWs) independently of the radio mechanism if the orbital period corresponds to the pulsation period. As eccentricity should be erased before a binary compactifies to $P_{\text{orb}}\sim$ hours (Peters, 1964), emissions at frequencies of $f_{\rm GW}=2/P_{\text{orb}}\approx 0.6\,\text{mHz}\times\left({1\,\text{hr}}/{P_{\text{orb}}}\right)$ ought to dominate the spectrum in this scenario. Within the next decade, a network of space-based, GW interferometers will become operational, namely the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017) plus the two Chinese-led projects Taiji (Luo et al., 2020) and TianQin (Huang et al., 2020). We focus on LISA and Taiji in this work, as they will be most sensitive in the $\gtrsim$ mHz band and are best-suited for detecting signals from known LPTs. We show that future GW observations can be leveraged to effectively confirm or rule out binary scenarios in a number of systems where electromagnetic data remain inconclusive.
@@
This paper is organized as follows. Multiwavelength observations and binary scenarios are reviewed in Section 2, to pave the way for estimates of the characteristic strains from LPTs in Section 3. Signal-to-noise ratios (S/Ns) are quantified in Section 3.1 for LISA alone and for a joint network in Section 3.2, where it is shown that several LPTs are promising GW detection candidates under favorable conditions. An outlook for future surveys is provided in Section 4. Astrophysical implications are also discussed, with conclusions presented in Section 5.
@@ 2 Constraints on binary nature
In this work, we pose the question: if LPTs (i) consist of a binary system involving at least one compact object and (ii) pulsate at the orbital period, would they be detectable by space-based GW interferometers? Given that the period is known in these systems, the main additional parameters needed for this assessment are the distance, $d$, and the chirp mass,
$\mathcal{M}=\frac{\left(M_{\star}M_{\rm c}\right)^{3/5}}{\left(M_{\star}+M_{\rm c}\right)^{1/5}}~,$ (1)
for primary and companion masses $M_{\star}$ and $M_{\rm c}$, respectively. While LPT distances are constrained to within a factor $\sim 2$ by their dispersion measures (DMs; cf. Price, Flynn & Deller, 2021), system masses (or indeed binary nature) are much more uncertain in most cases. In order to estimate the viable chirp-mass range and the viability of the hypothesis generally, we begin by reviewing the main constraints in this context.
@@ 2.1 White-dwarf plus M-dwarf
Two LPTs—GLEAM-X J0740–34 and ILT J1101+5521—have been confirmed as WDMDs with $P=P_{\rm orb}$ from broadband photometry, optical spectroscopy, and other channels (Hurley-Walker et al., 2024; de Ruiter et al., 2025; Rodriguez, 2025). Their data respectively indicate $M_{\rm c}\approx 0.19M_{\odot}$ and $M_{\rm c}\approx 0.32M_{\odot}$ (though cf. Rodriguez, 2025, for a Markov chain Monte Carlo parameter exploration). While primary masses cannot be dynamically-inferred with certainty due to the unknown inclination, the discovery papers suggest $M_{\star}\sim 0.6M_{\odot}$ and $M_{\star}\sim 0.8M_{\odot}$, respectively.
For ILT J1101+5521 however, de Ruiter et al. (2025) note that arguments based on Roche-lobe overflow indicate the tighter bound [^1] This restriction applies for an M-dwarf radius of $R_{\rm MD}=0.217R_{\odot}$. If the radius was lower by even a few percent ($R_{\rm WD}\lesssim 0.21R_{\odot}$), the primary mass becomes practically unconstrained (see their extended data figure 8). $M_{\star}\lesssim 0.3M_{\odot}$. In particular, observations of M dwarfs indicate a mass-radius relationship of $M_{\rm MD}/M_{\odot}\approx R_{\rm MD}/R_{\odot}$ (Parsons et al., 2018), with the lightest (confirmed) object being EBLM J0555–57AB with $M_{\rm MD}=0.084(4)M_{\odot}$ (von Boetticher et al., 2019). Using the Eggleton (1983) formula and assuming a Keplerian orbit, the radius of the Roche sphere reads
@@
[…]
for orbital separation $a$ and mass-ratio $q$. Since $R_{\rm MD}\leq R_{\rm R}$ is required for stability, an effective upper limit to $M_{\star}$ via $q$ is set by expression (2) if $P=P_{\rm orb}$. Such an argument further implies that LPTs with $P\lesssim 30$ min are unlikely to be WDMDs pulsing at the orbital period (see Table 1 for our considered sample [^2] Note that no distance estimate is provided by Anumarlapudi et al. (2025) for ASKAP J1448. They consider a plausible lower limit of $\sim 200$ pc, adopted here, with a maximum of $\sim 10$ kpc.). We exclude a WDMD scenario for J0630, GX J1627, and GPM J1839 for this reason (see also Sec. 2.3.1). For GX J1627 in particular, dedicated spectrophotometric follow-up practically rules out an M dwarf (Lyman et al., 2025). Alternative scenarios with either more compact progenitors, $P\neq P_{\rm orb}$, or isolated stars are required in these cases.
Table 1: Pulsation periods (in descending order) and dispersion-measure distances for the LPTs considered in this work. Full source names, and abbreviations used throughout, are provided in the notes below. For ILT/CHIME J1634+44, two orbital periods have been considered in the literature owing to interpulse/offset phenomena (see Sec. 2.3.1).
@@
[…]
With respect to the possibility that $P\neq P_{\rm orb}$, there are two known white-dwarf pulsars (J1912 and Ar Sco) with periods that do not correspond to the orbit but rather the spin of the primary (Marsh et al., 2016; Pelisoli et al., 2023, 2024). However, the radio profiles of these objects are notably different from the two confirmed WDMD LPTs. Their observed (highest peak) radio luminosities, $\sim 2\times 10^{26}$ erg/s at $\approx 1$ GHz, are much lower than the values of ILT J1101 ($\sim 1.5\times 10^{28}$ erg/s at $\approx$1 GHz) and GX J0740 ($\sim 10^{28}$ erg/s at $\approx$100 MHz); see appendix B.3 of Rodriguez (2025). Additionally, the degree of linear polarization in J1912 and Ar Sco are at the level of a few percent while the LPTs ILT J1101 and GX J0740 display values of at least $\approx 20\%$ (see table 1 in Qu & Zhang, 2025). Ar Sco and J1912 also shine persistently in X-rays, while LPTs do not (with the exception of A J1448; Anumarlapudi et al., 2025). The operative pulsation mechanisms and binary characteristics are therefore likely different between these astronomical classes; for this reason, we exclude J1912 and AR Sco from our sample (though the latter may be marginally visible for a multi-year folding-time given its distance of $\sim 110$ pc, chirp mass $\sim 0.5M_{\odot}$, and orbital period $\approx 3.57$ hours; Marsh et al., 2016). We also exclude ASKAP J183950.5–075635.0 (A J1839) from our sample as this source displays prominent interpulses at (main-pulse) phases of $\approx 0.5$ (Lee et al., 2025). While this cannot conclusively rule-out a binary origin for the source (cf. Bloot et al., 2025), these extra, half-period pulses are more easily explained by an isolated object emitting from both magnetic poles (see Suvorov, Dehman & Pons, 2025, for a magnetar model).
@@ 2.2 White dwarfs with compact companions
If the shortest period LPTs are not WDMDs, what type of systems could they be? While isolated neutron stars are viable – and perhaps even favored in most cases (Beniamini, Wadiasingh & Metzger, 2020; Cooper & Wadiasingh, 2024) – double degenerates are harder to rule out owing to their intrinsic dimness (see Lyman et al., 2025, for a discussion).
@@ 2.2.1 CHIME J0630+25
$\dot{P}_{\rm GW}$ $=-\frac{96G^{5/3}(4\pi^{2})^{5/3}}{5c^{5}}P_{\rm orb}^{-5/3}\mathcal{M}^{5/3}$ (5)
[…]
for Newton’s constant $G$ and speed of light $c$. Secondly, the lack of an obvious optical counterpart is concerning owing to the closeness of the source ($d\sim 100$ pc). Notably, a counterpart should be fainter than 22 magnitudes based on nondetections in Pan-STARRS (Kaiser et al., 2002), and many WDWDs are brighter. For example, ZTF J153932.16+502738.8 (J1539) has an orbital period of $\sim 7$ min and is optically visible despite its $\sim 2.3$ kpc distance (Burdge et al., 2019). Dust obscuration or binary inclination could alleviate this tension to some extent, as could additional optical monitoring given the relatively poor localization of the source, though these factors still challenge a WDWD interpretation. On the other hand, the fact that the best-fit $\dot{P}$ is negative hints at a binary, as emphasized by Dong et al. (2025a). Moreover, J1539 exhibits orbital decay that is consistent with equation (5), providing evidence for the viability of a double-compact case (Burdge et al., 2019).
All considered, we fix $\mathcal{M}_{\textrm{J}0630}=0.1M_{\odot}$ to avoid tension with expression (5), remarking that orbital decay could be accelerated by tides or electromagnetic interactions (decreasing the permitted $\mathcal{M}$) or decelerated if dissipative torques transfer angular momentum from the primary to the orbit (increasing it; see Marsh, Nelemans & Steeghs, 2004). We emphasize that this source, together with A J1839 and DA J1832 (see Lee et al., 2025; Wang et al., 2025), is probably the least likely candidate for a binary LPT. Nevertheless, our key point is that LISA can constrain binary scenarios in ways electromagnetic data cannot: even if $\mathcal{M}$ was lower by an order of magnitude, the closeness of this source together with its short period would permit detection if $P=P_{\rm orb}$ (Sec. 3.1).
@@ 2.2.2 GX J1627 and GPM J1839
GPM J1839 stands out amongst LPTs as this source has displayed a high duty cycle ($\sim 25\%$) going back to 1988 (Hurley-Walker et al., 2023). In a neutron-star scenario, this points towards aggressive crustal activity despite being old and ‘cold’ (Cooper & Wadiasingh, 2024), which could be used to implicitly constrain magnetar evolution (e.g. Dehman et al., 2020). This is supported by observations of linear-to-circular polarization conversion taking place in a manner remarkably reminiscent of the transient magnetar XTE J1810–197 (Men et al., 2025). If instead confirmed as a binary, independently of whether $P=P_{\rm orb}$, the fact that the radio luminosity exceeds that of AR Sco by several orders of magnitude indicates that pulsations from binaries involving dwarfs could be highly variable, important to account for in future radio surveys. Dedicated X-ray monitoring would also help constrain the nature of the source, as the current upper limit of $L^{\rm max}_{\rm X}\sim 10^{33}$ erg/s is relatively weak.
@@ 2.3 Neutron stars plus one
Except where the above exclusions apply, the chirp mass could be considerably higher if invoking a compact neutron-star/white-dwarf binary like PSR J1141–6545 ($P_{\rm orb}\approx 4.7$ hr and $\mathcal{M}\sim 1\,M_{\odot}$; Bailes et al., 2003) or a neutron/helium star binary like PSR J1928+1815 ($P_{\rm orb}\approx 3.6$ hr and $\mathcal{M}\sim 1.2M_{\odot}$; Yang et al., 2025). Binaries involving mature neutron stars are particularly difficult to constrain with multiwavelength data, much like isolated objects. The pulsation periods may correspond to the primary’s rotation rate in such cases – and indeed it has been argued that mass-loaded winds, magnetospheric twists, or interactions with a fallback disk could allow magnetars to reach ultra-long periods (Beniamini, Wadiasingh & Metzger, 2020; Beniamini et al., 2023; Suvorov, Dehman & Pons, 2025) – but a binary involving at least one neutron star is viable for many LPTs (e.g. Lyman et al., 2025). The final (most optimistic) scenario we consider corresponds to a J1141-like chirp mass of
@@
where the subscript indicates multiple possibilities for a compact companion (helium dwarf, white dwarf, or neutron star). For example, Turolla, Possenti & Treves (2005) suggested that GCRT J1745–3009 could contain a double neutron-star system with $\mathcal{M}\approx 1.13M_{\odot}$ and $P=P_{\rm orb}$ where radio emission originates from shocks within the interaction zone of the primary’s wind and the companion’s magnetosphere.
2.3.1 CHIME/ILT J1634+44
This source exhibits the unusual property that its emission is almost 100% circularly polarized (Dong et al., 2025b), although some pulses are nearly completely linearly polarized (Bloot et al., 2025). This contrasts with other LPTs and further emphasizes the likelihood of (at least) two distinct progenitor categories. An important feature of J1634 is the confident detection by CHIME of a negative period derivative, $\dot{P}\approx-9.03(11)\times 10^{-12}\text{ ss}^{-1}$ (Dong et al., 2025b). Despite the absence of any multiwavelength counterparts (with an X-ray limit of $L^{\rm max}_{\rm X}\lesssim 10^{32}$ erg/s), this provides strong evidence for a binary.
While the observed pulsation period $P=841.25$ s may correspond to the orbit, the existence of occasional, interpulse-like peaks with period $\sim 4206$ s could be related to a spin-orbit resonance (Bloot et al., 2025). In particular, these secondary ‘bursts’ were observed only at certain phases with different Stokes offsets depending on the observational epoch. The modulation period of $P=2103.1$ s was interpreted by Bloot et al. (2025) to be the orbital period assuming a 5:2 spin-orbit resonance.
Since both Bloot et al. (2025) and Dong et al. (2025b) mention that the radio data alone cannot strictly exclude an orbital period of $P=841.25$ s or a larger modulation period, we consider two scenarios for this source, with either $P_{\rm orb}=841.2$ s or $P_{\rm orb}=2103.1$ s. The period derivative quoted above then translates into chirp masses of $\mathcal{M}\approx 0.36M_{\odot}$ or $\mathcal{M}\approx 0.91M_{\odot}$ through expression (5), respectively, assuming GW emissions dominate angular momentum losses. The former aligns better with a WDMD or WDWD, while the latter is more consistent with a neutron star ‘companion’. We consider multiple scenarios for the $P=2103.1$ s case, since a unipolar inductor or tides (for instance) could accelerate inspiral and thus reduce the inferred $\mathcal{M}$ (see also Sec. 4).
We close this section by stressing that we have not attempted to present a complete survey of LPT scenarios, as these can be found in the respective discovery papers and literature (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). Indeed, our goal here concerns GW visibility, for which the exact binary makeup is unimportant (though is, of course, important for the implications of detections) Even if certain types of binarity may be disfavored in individual cases, we consider a range of chirp masses to estimate S/Ns. Larger or smaller $\mathcal{M}$ values can be easily substituted using the estimates provided in the following sections.